# Unveiling the Hidden Structure of Self-Attention via Kernel Principal Component Analysis

### v1

@@ Abstract
The remarkable success of transformers in sequence modeling tasks, spanning various applications in natural language processing and computer vision, is attributed to the critical role of self-attention. Similar to the development of most deep learning models, the construction of these attention mechanisms rely on heuristics and experience. In our work, we derive self-attention from kernel principal component analysis (kernel PCA) and show that self-attention projects its query vectors onto the principal component axes of its key matrix in a feature space. We then formulate the exact formula for the value matrix in self-attention, theoretically and empirically demonstrating that this value matrix captures the eigenvectors of the Gram matrix of the key vectors in self-attention. Leveraging our kernel PCA framework, we propose Attention with Robust Principal Components (RPC-Attention), a novel class of robust attention that is resilient to data contamination. We empirically demonstrate the advantages of RPC-Attention over softmax attention on the ImageNet-1K object classification, WikiText-103 language modeling, and ADE20K image segmentation task.
@@ 1 Introduction
Transformers [68] have emerged as the preeminent model for tackling a myriad of challenging problems in natural language processing [3, 16, 1, 51, 7, 33, 13], computer vision [18, 64, 37, 48, 32, 20], reinforcement learning [10, 29, 78, 34], and other applications [77, 53, 30, 73]. The effectiveness of transformers is rooted in their ability to learn from unlabeled data and take advantage of pre-trained models for downstream tasks that involve diverse data modalities with limited supervision [49, 50, 74, 81, 52]. At the core of the transformer’s success lies the self-attention mechanism, which serves as the fundamental building block of a transformer model. This mechanism enables each token in a sequence to aggregate information from other tokens by computing a weighted average based on similarity scores between their feature representations. Facilitating dynamic interactions among tokens, this attention mechanism allows tokens to selectively attend to others, thereby attaining a contextual representation [4, 45, 36]. The flexibility in capturing diverse syntactic and semantic relationships is an important factor contributing to the success of transformers [62, 70, 27, 69].
@@ 2 Principal Component Analysis of Attention
The covariance matrix in the feature space ${\bm{C}}$ is given by
${\bm{C}}=\frac{1}{N}\sum_{j=1}^{N}\tilde{{\bm{\varphi}}}({\bm{k}}_{j})\tilde{{\bm{\varphi}}}({\bm{k}}_{j})^{\top}$
and its eigenvector expansion is
${\bm{C}}{\bm{u}}_{d}=\lambda_{d}{\bm{u}}_{d},\,\,\,d=1,\dots,D_{v}.$ (5)
@@ 2.1 Deriving Attention from Kernel PCA
${\bm{h}}_{i}(d)$ $={\bm{\varphi}}({\bm{q}}_{i})^{\top}{\bm{u}}_{d}$
$=\sum_{j=1}^{N}a_{dj}{\bm{\varphi}}({\bm{q}}_{i})^{\top}\tilde{{\bm{\varphi}}}({\bm{k}}_{j})$
$=\sum_{j=1}^{N}a_{dj}{\bm{\varphi}}({\bm{q}}_{i})^{\top}({\bm{\varphi}}({\bm{k}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}{\bm{\varphi}}({\bm{k}}_{j^{\prime}}))$
@@
$=\sum_{j=1}^{N}\frac{k({\bm{q}}_{i},{\bm{k}}_{j})}{g({\bm{q}}_{i})}\Big{(}\frac{a_{dj}}{g({\bm{k}}_{j})}-\frac{1}{N}\sum_{j^{\prime}=1}^{N}\frac{a_{dj^{\prime}}}{g({\bm{k}}_{j})}\Big{)}$
$=\sum_{j=1}^{N}\frac{k({\bm{q}}_{i},{\bm{k}}_{j})}{g({\bm{q}}_{i})}v_{dj},$ (8)
@@ 2.2.1 Projection Error Minimization
In the derivation above, we leverage the orthonormality of $\{{\bm{u}}_{d}\}_{d=1}^{D_{v}}$ and $h_{di}={\bm{\varphi}}({\bm{q}}_{i})^{\top}{\bm{u}}_{d}$. Here, notice that we can compute $\|{\bm{\varphi}}({\bm{q}}_{i})\|^{2}$ from ${\bm{q}}_{i}$ and $\{{\bm{k}}_{j}\}_{j=1}^{N}$ as
$\|{\bm{\varphi}}({\bm{q}}_{i})\|^{2}=\exp({\bm{q}}_{i}^{\top}{\bm{q}}_{i}/\sqrt{D})/(\sum_{j=1}^{N}\exp({\bm{q}}_{i}^{\top}{\bm{k}}_{j}/\sqrt{D}))^{2}.$
In Fig. 1, we empirically show that a transformer model minimizes the projection loss $J_{\text{proj}}$ during training. Here, we train a vision transformer [18], ViT-tiny model in particular, on the ImageNet-1K object classification task and compute the average of $J_{\text{proj}}$ across attention heads and layers. This result suggests that during training, the transformer learns to perform kernel PCA at each self-attention layer by implicitly minimizing the projection loss $J_{\text{proj}}$.
@@
Thus, the value vector ${\bm{v}}_{j}$, $j=1,\dots,N$, in self-attention layers converge to the values specified in Theorem 1. We provide more details on the computation of $J_{\text{proj}}$ in Appendix C.
2.2.2 Learning the Eigenvectors of the Gram Matrix
@@
We empirically verify that $\frac{\widetilde{{\bm{K}}}_{{\bm{\varphi}}}\hat{{\bm{a}}}_{d}}{N\hat{{\bm{a}}}_{d}}={\bm{\gamma}}=[\gamma_{1},\dots,\gamma_{N}]$ where $\gamma_{1}=\dots=\gamma_{N}=\text{const}$, which confirms that $\hat{{\bm{a}}}_{d}$ is an eigenvector of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$. In particular, in Fig. 2(Left), we plot the average pairwise absolute differences of $\gamma_{i}$ and $\gamma_{j}$, $i\neq j$, $i,j=1,\dots,N$, for each principal component axis of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$. The results are averaged over all attention heads and all layers in the 12-layer ViT-tiny model trained on ImageNet-1K. As can be seen in our figure, the absolute difference between any pair of $\gamma_{i}$ and $\gamma_{j}$ is almost 0 with a very small standard deviation. Similar results are observed at each layer when averaging over all attention heads in that layer. In Fig. 2(Right), we show the results for Layers 2, 5, 8, and 11 in the model and a plot of all layers is visualized in Appendix E.2. For comparison, we observe that the max, min, mean, and median of the absolute values of these $D_{v}$ eigenvalues, averaged over all attention heads and layers, are 648.46, 4.65, 40.07, and 17.73, respectively, which are much greater than the values of $|\gamma_{i}-\gamma_{j}|$. These results empirically justify that $\frac{\widetilde{{\bm{K}}}_{{\bm{\varphi}}}\hat{{\bm{a}}}_{d}}{N\hat{{\bm{a}}}_{d}}=\text{const}$ and the value matrix ${\bm{V}}$ captures the eigenvectors of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$ after the transformer model is trained, as suggested in Eqn. (11).
@@ 3.2 Attention with Robust Principal Components
$\arg\min_{{\bm{S}}}\mathcal{L}({\bm{L}},{\bm{S}},{\bm{Y}})$ $=\mathcal{S_{\lambda/\mu}}({\bm{K}}-{\bm{L}}+\mu^{-1}{\bm{Y}})$
$\arg\min_{{\bm{L}}}\mathcal{L}({\bm{L}},{\bm{S}},{\bm{Y}})$ $=\mathcal{D}_{\mu}({\bm{K}}-{\bm{S}}-\mu^{-1}{\bm{Y}}).$
$\mathcal{D_{\mu}}$ finds a low-rank approximation of ${\bm{K}}-{\bm{S}}-\mu^{-1}{\bm{Y}}$. Thus, we obtain an approximation of the above equation by replacing $\mathcal{D_{\mu}}$ by a low-rank approximation operator. Such an approximation
@@
takes a step towards the minimum of $\mathcal{L}({\bm{L}},{\bm{S}},{\bm{Y}})$ when fixing ${\bm{S}}$ and ${\bm{Y}}$. It has been empirically observed and theoretically proven that the output matrix ${\bm{H}}$ of self-attention is low-rank, a phenomenon known as over-smoothing or rank collapse [56, 71, 17]. Therefore, we can replace $\mathcal{D_{\mu}}$ by a self-attention operator. The ADMM method applied to self-attention, which we name Principal Attention Pursuit (PAP), is given by Algorithm 1. We define our RPC-Attention as
@@ 4 Experimental Results
Throughout our experiments, we compare the performance of our proposed models with the baseline softmax attention of the same configuration. All of our results are averaged over 5 runs with different seeds and run on 4 A100 GPU. Details on the models and training settings are provided in Appendix A and additional experimental results are provided in Appendix E. Primarily, we focus on a ViT-tiny model backbone [18], but included in the appendix are experiments on a larger model backbone, ViT-small, and a state of the art (SOTA) robust model, Fully Attentional Networks (FAN)  [80].
@@ 4.3 Validating the Benefits of Scaled Attention
In this section, we provide an empirical comparison between the Scaled Attention in Remark 3 and the softmax attention. We train an asymmetric ViT-tiny model with two different versions of Scaled Attention. While the exact value of ${\bm{S}}$ can be mathematically formulated as in Remark 3, it might lead to numerical errors that are difficult to handle. Therefore, in the first version of Scaled Attention, we let ${\bm{S}}$ be a learned parameter matrix in each layer (Scaled Attention $S$ in Fig. 3(right)). In the second version of Scaled Attention (Scaled Attention $\alpha\times\text{Asym}$ in Fig. 3(right)), given that $s_{jj^{\prime}}=g(k_{j^{\prime}})/g(k_{j})$, we rewrite ${\bm{S}}$ as the product of a symmetric softmax
@@
attention matrix and the reciprocal of its transpose. The model learns this reciprocal by a learnable scalar $\alpha$ and we let ${\bm{S}}=\alpha{\bm{A}}_{Sym}$, where ${\bm{A}}_{Sym}$ is the symmetric softmax attention. More details are in Appendix E.1.
@@ 5 Related Works
Theoretical Perspectives of Attention Mechanisms. The study of the attention mechanism in transformers through different theoretical frameworks has been expanding. [66] shows that attention can be analyzed as a kernel smoother over the inputs using an appropriate kernel score that is the similarities between them. [12, 72, 31] reduce the quadratic complexity of transformers by linearizing the softmax kernel to improve the computational and memory efficiency. In addition, there are works interpreting transformers using the frameworks of ordinary/partial differential equations [38, 54, 22, 21] and from probabilistic viewpoints with Gaussian Mixture Models [19, 60, 76]. [11] provides a new perspective by emphasizing the asymmetry of the softmax kernel and recovers the self-attention mechanism from an asymmetric Kernel Singular Value Decomposition (KSVD) using the duality of the optimization problem. Another related work views transformers from the perspective of Support Vector Machines [61]. We discuss [11] and [61] in more details in Appendix F. Separate from these works, our kernel PCA perspective derives softmax attention as a projection of the query vectors in a feature space. Using our framework, we are able to predict the exact explicit form of the value matrix in self-attention, demonstrating that this matrix captures the eigenvectors of the Gram matrix of the key vectors in a feature space. Our work is the first to show this insight.
Robustness of Transformers. There have been many works studying the robustness of Vision Transformers (ViT) [18] against different types of attacks [5, 46, 58, 80]. Recent work that serves to address this include [40], whereby new training strategies and architectural adjustments are proposed to improve the robustness of ViT. Our RPC-Attention is orthogonal to these methods.
@@ A.1 Robustness against Data Corruptions
On ImageNet-1K, ImageNet-C, ImageNet-A and ImageNet-R, we report the top-1 accuracies for all experiments. We include top-5 accuracies on ImageNet-1K. On ImageNet-C, the standard metric for evaluation is the mean Corruption Error (mCE). To calculate this, we average the top-1 error rate for each corruption type across the 5 levels of severity and divide them by AlexNet’s average errors, then take the final average across all corruption types. We report the area under the precision-recall curve (AUPR) for ImageNet-O which requires anomaly scores. The score is obtained by taking the negative of the highest softmax probability output by the model. The direction of increasing or decreasing values of these metrics signifying greater robustness will be indicated in the table with an arrow.

### v2

@@ Abstract
The remarkable success of transformers in sequence modeling tasks, spanning various applications in natural language processing and computer vision, is attributed to the critical role of self-attention. Similar to the development of most deep learning models, the construction of these attention mechanisms relies on heuristics and experience. In our work, we derive self-attention from kernel principal component analysis (kernel PCA) and show that self-attention projects its query vectors onto the principal component axes of its key matrix in a feature space. We then formulate the exact formula for the value matrix in self-attention, theoretically and empirically demonstrating that this value matrix captures the eigenvectors of the Gram matrix of the key vectors in self-attention. Leveraging our kernel PCA framework, we propose Attention with Robust Principal Components (RPC-Attention), a novel class of robust attention that is resilient to data contamination. We empirically demonstrate the advantages of RPC-Attention over softmax attention on the ImageNet-1K object classification, WikiText-103 language modeling, and ADE20K image segmentation task. The code is publicly available at https://github.com/rachtsy/KPCA_code.
@@ 1 Introduction
Transformers [83] have emerged as the preeminent model for tackling a myriad of challenging problems in natural language processing [4, 20, 1, 66, 9, 39, 17], computer vision [22, 78, 43, 63, 38, 24], reinforcement learning [13, 35, 94, 40], and other applications [93, 68, 36, 88]. The effectiveness of transformers is rooted in their ability to learn from unlabeled data and take advantage of pre-trained models for downstream tasks that involve diverse data modalities with limited supervision [64, 65, 90, 97, 67]. At the core of the transformer’s success lies the self-attention mechanism, which serves as the fundamental building block of a transformer model. This mechanism enables each token in a sequence to aggregate information from other tokens by computing a weighted average based on similarity scores between their feature representations. Facilitating dynamic interactions among tokens, this attention mechanism allows tokens to selectively attend to others, thereby attaining a contextual representation [5, 56, 60, 42]. The flexibility in capturing diverse syntactic and semantic relationships is an important factor contributing to the success of transformers [77, 85, 33, 84, 53].Please correspond to: tanmn@nus.edu.sg
@@ 2 Principal Component Analysis of Attention
Letting ${\bm{C}}$ be the covariance matrix of the centered data in feature space, its eigenvector expansion is
${\bm{C}}=\frac{1}{N}\sum_{j=1}^{N}\tilde{{\bm{\varphi}}}({\bm{k}}_{j})\tilde{{\bm{\varphi}}}({\bm{k}}_{j})^{\top};\quad{\bm{C}}{\bm{u}}_{d}=\lambda_{d}{\bm{u}}_{d},\,\,\,d=1,\dots,D_{v}.$ (5)
@@ 2.1 Deriving Attention from Kernel PCA
${\bm{h}}_{i}(d)$ $={\bm{\varphi}}({\bm{q}}_{i})^{\top}{\bm{u}}_{d}=\sum_{j=1}^{N}a_{dj}{\bm{\varphi}}({\bm{q}}_{i})^{\top}\tilde{{\bm{\varphi}}}({\bm{k}}_{j})=\sum_{j=1}^{N}a_{dj}{\bm{\varphi}}({\bm{q}}_{i})^{\top}({\bm{\varphi}}({\bm{k}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}{\bm{\varphi}}({\bm{k}}_{j^{\prime}}))$
@@
$=\sum_{j=1}^{N}\frac{k({\bm{q}}_{i},{\bm{k}}_{j})}{g({\bm{q}}_{i})}\Big{(}\frac{a_{dj}}{g({\bm{k}}_{j})}-\frac{1}{N}\sum_{j^{\prime}=1}^{N}\frac{a_{dj^{\prime}}}{g({\bm{k}}_{j})}\Big{)}=\sum_{j=1}^{N}\frac{k({\bm{q}}_{i},{\bm{k}}_{j})}{g({\bm{q}}_{i})}v_{dj},$ (8)
@@ 2.2.1 Projection Error Minimization
In the derivation above, we leverage the orthonormality of $\{{\bm{u}}_{d}\}_{d=1}^{D_{v}}$ and $h_{di}={\bm{\varphi}}({\bm{q}}_{i})^{\top}{\bm{u}}_{d}$. Here, notice that we can compute $\|{\bm{\varphi}}({\bm{q}}_{i})\|^{2}$ from ${\bm{q}}_{i}$ and $\{{\bm{k}}_{j}\}_{j=1}^{N}$ as $\|{\bm{\varphi}}({\bm{q}}_{i})\|^{2}=\exp({\bm{q}}_{i}^{\top}{\bm{q}}_{i}/\sqrt{D})/(\sum_{j=1}^{N}\exp({\bm{q}}_{i}^{\top}{\bm{k}}_{j}/\sqrt{D}))^{2}$. In Fig. 1, we empirically show that a transformer model minimizes the projection loss $J_{\text{proj}}$ during training. Here, we train a vision transformer [22], ViT-tiny model in particular, on the ImageNet-1K classification task and compute the average of $J_{\text{proj}}$ across attention heads and layers. This result suggests that during training, the transformer learns to perform PCA at each self-attention layer by implicitly minimizing the projection loss $J_{\text{proj}}$. Thus, the value vector ${\bm{v}}_{j}$, $j=1,\dots,N$, in self-attention layers converge to the values specified in Theorem 1. We provide more details on the computation of $J_{\text{proj}}$ in Appendix C.
2.2.2 Learning Eigenvectors of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$ in Eqn. (11)
@@
We empirically verify that $\frac{\widetilde{{\bm{K}}}_{{\bm{\varphi}}}\hat{{\bm{a}}}_{d}}{N\hat{{\bm{a}}}_{d}}={\bm{\gamma}}=[\gamma_{1},\dots,\gamma_{N}]$ where $\gamma_{1}=\dots=\gamma_{N}=\text{const}$, which confirms that $\hat{{\bm{a}}}_{d}$ is an eigenvector of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$. In particular, in Fig. 2(Left), we plot the average pairwise absolute differences of $\gamma_{i}$ and $\gamma_{j}$, $i\neq j$, $i,j=1,\dots,N$, for each principal component axis of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$. The results are averaged over all attention heads and all layers in the 12-layer ViT-tiny model trained on ImageNet-1K. As can be seen in our figure, the absolute difference between any pair of $\gamma_{i}$ and $\gamma_{j}$ is almost 0 with a very small standard deviation. Similar results are observed at each layer when averaging over all attention heads in that layer. In Fig. 2(Right), we show the results for Layers 2, 5, 8, and 11 in the model. For comparison, we observe that the max, min, mean, and median of the absolute values of these $D_{v}$ eigenvalues, averaged over all attention heads and layers, are 648.46, 4.65, 40.07, and 17.73, respectively, which are much greater than the values of $|\gamma_{i}-\gamma_{j}|$. These results empirically justify that $\frac{\widetilde{{\bm{K}}}_{{\bm{\varphi}}}\hat{{\bm{a}}}_{d}}{N\hat{{\bm{a}}}_{d}}=\text{const}$ and the value matrix ${\bm{V}}$ captures the eigenvectors of $\widetilde{{\bm{K}}}_{{\bm{\varphi}}}$ after the transformer model is trained, as suggested in Eqn. (11).
@@ 3.2 Attention with Robust Principal Components
$\arg\min_{{\bm{S}}}\mathcal{L}({\bm{L}},{\bm{S}},{\bm{Y}})=\mathcal{S_{\lambda/\mu}}({\bm{K}}-{\bm{L}}+\mu^{-1}{\bm{Y}});\,\arg\min_{{\bm{L}}}\mathcal{L}({\bm{L}},{\bm{S}},{\bm{Y}})=\mathcal{D}_{\mu}({\bm{K}}-{\bm{S}}-\mu^{-1}{\bm{Y}}).$
$\mathcal{D_{\mu}}$ finds a low-rank approximation of ${\bm{K}}-{\bm{S}}-\mu^{-1}{\bm{Y}}$. Thus, we obtain an approximation of the above equation by replacing $\mathcal{D_{\mu}}$ by a low-rank approximation operator. Such an approximation takes a
@@
step towards the minimum of $\mathcal{L}({\bm{L}},{\bm{S}},{\bm{Y}})$ when fixing ${\bm{S}}$ and ${\bm{Y}}$. It has been empirically observed and theoretically proven that the output matrix ${\bm{H}}$ of self-attention is low-rank, a phenomenon known as over-smoothing or rank collapse [71, 86, 21]. Therefore, we can replace $\mathcal{D_{\mu}}$ by a self-attention operator. The ADMM method applied to self-attention, which we name Principal Attention Pursuit (PAP), is given by Algorithm 1. We define our RPC-Attention as
@@ 4 Experimental Results
Throughout our experiments, we compare the performance of our proposed models with the baseline softmax attention of the same configuration. All of our results are averaged over 5 runs with different seeds and run on 4 A100 GPU. Details on the models and training settings are provided in Appendix A and additional experimental results are provided in Appendix E. Primarily, we focus on a ViT-tiny model backbone [22], but included in the appendix are experiments on a larger model backbone, ViT-base, and a state of the art (SOTA) robust model, Fully Attentional Networks (FAN)  [96].
@@ 4.3 Validating the Benefits of Scaled Attention
In this section, we provide an empirical comparison between the Scaled Attention in Remark 3 and the softmax attention. We train an asymmetric ViT-tiny model with two different versions of Scaled Attention. While the exact value of ${\bm{S}}$ can be mathematically formulated as in Remark 3, it might lead to numerical errors that are difficult to handle. Therefore, in the first version of Scaled Attention, we let ${\bm{S}}$ be a learned parameter matrix in each layer (Scaled Attention $S$ in Fig. 3(right)). In the second version of Scaled Attention (Scaled Attention $\alpha\times\text{Asym}$ in Fig. 3(right)), given that $s_{jj^{\prime}}=g(k_{j^{\prime}})/g(k_{j})$, we rewrite ${\bm{S}}$ as the product of a symmetric softmax attention matrix and the reciprocal of its transpose. The model learns this reciprocal by a learnable scalar $\alpha$ and we let ${\bm{S}}=\alpha{\bm{A}}_{Sym}$, where ${\bm{A}}_{Sym}$ is the symmetric softmax attention. More details are in Appendix E.1.
@@ 5 Related Works
Theoretical Perspectives of Attention Mechanisms. The study of the attention mechanism in transformers through different theoretical frameworks has been expanding. [81] shows that attention can be analyzed as a kernel smoother over the inputs using an appropriate kernel score that is the similarities between them. [16, 87, 37, 55] reduce the quadratic complexity of transformers by linearizing the softmax kernel to improve the computational and memory efficiency. In addition, there are works interpreting transformers using the frameworks of ordinary/partial differential equations [50, 51, 44, 69, 27, 26] and from probabilistic viewpoints with Gaussian Mixture Models [52, 23, 75, 54, 92]. [15] provides a new perspective by emphasizing the asymmetry of the softmax kernel and recovers the self-attention mechanism from an asymmetric Kernel Singular Value Decomposition (KSVD) using the duality of the optimization problem. Another related work views transformers from the perspective of Support Vector Machines [57, 76]. We discuss [15] and [76], as well as the approaches that use matrix decomposition and iterative algorithms in deep models, in Appendix F. Separate from these works, our kernel PCA perspective derives softmax attention as a projection of the query vectors in a feature space. Using our framework, we are able to predict the exact explicit form of the value matrix in self-attention, demonstrating that this matrix captures the eigenvectors of the Gram matrix of the key vectors in a feature space. Our work is the first to show this insight.
Robustness of Transformers. There have been many works studying the robustness of Vision Transformers (ViT) [22] against different types of attacks [7, 61, 73, 96]. Recent work that serves to address this include [46], whereby new training strategies and architectural adjustments are proposed to improve the robustness of ViT. In addition, [58] suggests employing a Mahalanobis distance metric to calculate attention weights, expanding the feature space along directions of high contextual relevance, thereby enhancing the model’s robustness. Also, [29] adapts traditional robust kernel density estimation techniques to create new classes of transformers that are resilient to adversarial attacks and data contamination. [14, 10] integrate a Gaussian process into attention for out-of-distribution detection, and [80] develops equivariant neural functional networks for transformers. Our RPC-Attention is orthogonal to these methods.
@@ 6 Concluding Remarks
Acknowledgments and Disclosure of Funding
This research / project is supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-012-SGIL). This research / project is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (A-8002040-00-00, A-8002039-00-00). This research / project is also supported by the NUS Presidential Young Professorship Award (A-0009807-01-00).
@@ A.1 Robustness against Data Corruptions
On ImageNet-1K, ImageNet-A and ImageNet-R, we report the top-1 accuracies for all experiments. We include top-5 accuracies on ImageNet-1K. On ImageNet-C, the standard metric for evaluation is the mean Corruption Error (mCE). To calculate this, we average the top-1 error rate for each corruption type across the 5 levels of severity and divide them by AlexNet’s average errors, then take the final average across all corruption types. We report the area under the precision-recall curve (AUPR) for ImageNet-O which requires anomaly scores. The score is obtained by taking the negative of the highest softmax probability output by the model. The direction of increasing or decreasing values of these metrics signifying greater robustness will be indicated in the table with an arrow.